### v1

@@ 9 Conclusion
This paper introduces a simple idea: whereas information is valuable because it leads to better decisions, conceptual knowledge is valuable because it leads to better information. We formalize this idea and study its consequences. Conceptual knowledge is more valuable when the unknown states in an environment are more reducible: when they can be represented as lower-dimensional combinations of high-dimensional features. Its value is non-monotone in the quantity and quality of available data, and vanishes with infinite data. Deeper knowledge is (weakly) more valuable and allows one to attain the same welfare with less data. This trade-off between knowledge and data highlights the importance of interventions that give people knowledge as well as data. It also highlights an important role humans play in human-AI interactions: knowing how to “ask the right questions.”
Stemming from this paper are several avenues for future research. One is to analyze the trade-off between knowledge and data in a consumer choice setting. This would require specifying the “price” of conceptual knowledge vis-à-vis observations of data. Given a price schedule, one could ask many questions about knowledge and data: Are they complements or substitutes? Are they normal or inferior? How do these statuses depend on the states’ reducibility?
Another avenue is to consider competitive and persuasive forces. If there were many ways of interpreting the states’ structural relationships (i.e., many “models” of the agent’s environment), which interpretation would he choose? How could his choice be manipulated by a strategic communicator with competing incentives?
A third avenue is to make our framework dynamic. For example, the agent could take actions that generate outcomes observed by future agents. This would allow us to study how conceptual knowledge is “discovered” and passed down to new generations. This discovery process has been studied by other authors (Carnehl and Schneider,, 2025; Gans,, 2025); blending their models and insights with ours may bear fruit.
@@ Appendix A Additional material
This section connects our paper to the literatures on machine and statistical learning, which study how to derive predictive functions from data.333333 See Bishop, (2006) or Hastie et al., (2009) for textbook treatments. First, we show that our framework (described in Section 3) can be used to study Bayesian learning about real-valued functions. Second, we show how the agent’s prior derives from his “(conceptual) model” of an unknown function.
@@
The agent knows $\phi_{1},\ldots,\phi_{K}$ but not $\theta\equiv(\theta_{1},\ldots,\theta_{K})$, so learning about $f$ is equivalent to learning about $\theta$.343434 Assuming $\theta$ is normally distributed is equivalent to assuming $\{f(x)\}_{x\in\mathcal{X}}$ follows a Gaussian process. Such processes arise in the economic literature on learning and information acquisition—see, e.g., Bardhi, (2024), Davies, (2024), Ilut and Valchev, (2025), or Laajaj and Macours, (2024). See also Bishop, (2006, Section 6.4) or Rasmussen and Williams, (2006) for more information about Gaussian processes and their applications. Moreover, suppose the agent draws an input $x\in\mathcal{X}$ uniformly at random, and predicts the “output” $y\in\mathbb{R}$ with conditional distribution
@@
A1.2 Conceptual models
Suppose the agent knows about a collection $\psi_{1},\ldots,\psi_{J}\in\mathcal{F}$ of “attributes” that (partially) mediate the relationship between inputs and outputs.353535 This $J$ is the same as the depth parameter defined in Section 6. These attributes are linearly independent (but not necessary orthogonal) elements of the function space $\mathcal{F}$. They map inputs to known, measurable quantities.
The agent uses $\psi_{1},\ldots,\psi_{J}$ to build a “(conceptual) model” $m\in\mathcal{F}$ that approximates the unknown function $f$. This model is a linear combination of attributes: there is a(n unknown) vector $\beta\equiv(\beta_{1},\ldots,\beta_{J})\in\mathbb{R}^{J}$ such that
@@
of $m(x)$ with respect to $\psi_{k}(x)$ does not depend on the input $x$. In this way, the model $m$ captures the generalizable structure of $f$ that is common to all inputs. In contrast, the model’s “approximation error”
@@
of $\mathcal{F}$ spanned by the attributes. It corresponds to a subspace
@@
of the Euclidean space $\Theta\equiv\mathbb{R}^{K}$ containing the unknown coordinate vector $\theta$. Concretely, if $\theta^{j}\equiv(\theta_{1}^{j},\ldots,\theta_{K}^{j})$ contains the (known) coordinates of the $j^{\text{th}}$ attribute $\psi_{j}$ over the orthonormal basis $\mathcal{B}$, then
@@
and $1/2\leq t\leq 1/2$. Increasing $t$ from 1/2-1/2 to $1/2$ rotates $w^{(1)}$ clockwise from (1,0)(-1,0) to $(1,0)$. The value363636 We obtain (A4) by substituting $K=2$, the prior variance matrix (7), the covariate $w^{(1)}=(\sin(\pi t),\cos(\pi t))$, and the sample size $n=1$ into (A2).
@@
If $\mathcal{S}$} is representative, then its value $\pi({\mathcal{S}})$ does not fall when $\lambda_{1},\ldots,\lambda_{K}$ undergo a MPS.
@@
If the sample $\mathcal{S}$} is not representative, then its value can fall when the eigenvalues of $\Sigma$ undergo a MPS. This happens, for example, when ${\mathcal{S}}=\{(w^{(1)},y^{(1)})\}$ is a singleton and $w^{(1)}$ is an eigenvector of $\Sigma$ corresponding to an eigenvalue that falls under the MPS.373737 For example, if $\theta$ has prior variance (7), then the value of ${\mathcal{S}}=\{(w^{(1)},y^{(1)})\}$ is decreasing in $\rho$ when $w^{(1)}=v_{2}$. Such a sample is “non-spanning”: the rank
@@
of $G$. The agent cannot learn about these components from $\mathcal{S}$} directly. But he can learn about them indirectly if he knows how they relate structurally to the components that belong to $\mathrm{col}(G)$.
@@
for each $k\in\{1,\ldots,K\}$.383838 The matrix $\operatorname{\mathbb{V}}(\theta_{\mathcal{S}})$ is invertible because it is a leading principal submatrix of an invertible matrix. Then the posterior variance matrix has trace
@@
is the prior variance of $\theta_{k}$ left unexplained $\theta_{\mathcal{S}}$.393939 We derive (A5)–(A7) in Appendix B. The first two terms on the RHS of (A5) are “sampling errors” that depend on how much information $\mathcal{S}$} contains about the on-support components $\theta_{1},\ldots,\theta_{R}$. The third term is an “extrapolation error” that depends on how much information these components contain about the off-support components $\theta_{R+1},\ldots,\theta_{K}$. Whereas the sampling error can be reduced by collecting more (or less noisy) data on $\theta_{1},\ldots,\theta_{R}$, the extrapolation error cannot. It can only be reduced by knowing more about how the on- and off-support components of $\theta$ relate structurally.
@@
This example highlights the importance of conceptual knowledge when making out-of-sample predictions. If the agent did not know how the states related structurally (i.e., if $\rho=0$), then his extrapolation error could be arbitrarily large and the value of his sample could be arbitrarily small. Knowing how the states relate allows him to use data on the on-support components of $\theta$ to learn about the off-support components. This lowers his extrapolation error and ensures his sample has some value, even if the sampling frame is limited.
@@
    does not fall when $\lambda_{1},\ldots,\lambda_{K}$ undergo a MPS.
@@ Appendix B Proofs
for each $k\in\{1,\ldots,K\}$ gives $\tau_{k}^{\prime}\in[\tau_{k},\tau_{k+1})$ if and only if $K/\overline{\lambda}\in[\eta_{k},\eta_{k+1})$. But $\eta_{k}$ is non-increasing in $k$ because $\lambda_{k+1}\leq\lambda_{k}$ and therefore
@@
$\frac{\partial}{\partial\rho}\mathopen{}\mathclose{{\left[\pi^{*}\big{\rvert}_{\rho>\rho^{\prime}}}}\right]$ $=\frac{(K-1)\sigma^{2}}{K}\mathopen{}\mathclose{{\left(1+\mathopen{}\mathclose{{\left(\frac{1}{1+\tau\mathopen{}\mathclose{{\left(1+\rho(K+1)}}\right)}}}\right)^{2}}}\right)$
@@
    Now
@@
    Fix $n\geq 0$. Now $\pi^{(J)}$ is non-increasing in $J$ (by Lemma 5), so if $\pi^{(J)}\geq\pi_{0}$ then $\pi^{(J+1)}\geq\pi_{0}$. Thus
@@
$=\frac{1}{\sigma_{u}^{4}}\Sigma^{-1}G\Sigma^{-1}+\frac{2}{\sigma_{u}^{6}}G^{2}\Sigma^{-1}+\frac{1}{\sigma_{u}^{8}}G^{3},$
which is the sum of three matrices with strictly positive traces. Thus

### v2

@@ 9 Conclusion
This paper introduces a simple idea: whereas information is valuable because it leads to better decisions, conceptual knowledge is valuable because it leads to better information. We formalize this idea and study its consequences. Conceptual knowledge is more valuable when payoff-relevant unknowns are more “reducible”: when they can be explained with fewer common concepts. Its value is non-monotone in the quantity and quality of available data, and vanishes with infinite data. Deeper conceptual knowledge is (weakly) more valuable and compensates for having less data. So we can improve interventions that give people data by giving them conceptual tools to interpret data. We can also improve human-AI interactions by recognizing humans’ comparative advantage: knowing how to “ask the right questions.”
Stemming from this paper are several avenues for future research. One is to analyze the trade-off between concepts and data in a consumer choice setting. This would require specifying the “price” of acquiring concepts vis-à-vis observations of data. Given a price schedule, one could ask many questions about concepts and data: Are they complements or substitutes? Are they normal or inferior? How do these statuses depend on the states’ reducibility?
Another avenue is to make our framework dynamic. For example, the agent could take actions that generate outcomes observed by future agents. This would allow us to study how conceptual knowledge is “discovered” and passed down to new generations. This discovery process has been studied by other authors (Carnehl and Schneider,, 2025; Gans,, 2025); blending their models and insights with ours may bear fruit.
A third avenue is to study the role of conceptual knowledge in social learning and technology adoption. This paper and its empirical sibling (Sankar et al.,, 2025) came from studying agricultural settings, where heterogeneity in outcomes and the inability to generalize slows learning (Conley and Udry,, 2010; Munshi,, 2004; Tjernström,, 2017) and technology adoption (Alidaee,, 2023; BenYishay and Mobarak,, 2019; Laajaj and Macours,, 2025). These frictions could be removed by giving people conceptual tools that enable them to generalize. This paper offers a framework for studying such tools and their value.
@@ Appendix A Additional material
This section connects our paper to the literatures on machine and statistical learning, which study how to derive predictive functions from data. [^42] See Bishop, (2006) or Hastie et al., (2009) for textbook treatments. First, we show that our framework (described in Section 3) can be used to study Bayesian learning about real-valued functions. Second, we show how the agent’s prior derives from his approximating model of an unknown function.
@@
The agent knows $\phi_{1},\ldots,\phi_{K}$ but not $\theta\equiv(\theta_{1},\ldots,\theta_{K})$, so learning about $f$ is equivalent to learning about $\theta$. [^43] Assuming $\theta$ is normally distributed is equivalent to assuming $\{f(x)\}_{x\in\mathcal{X}}$ follows a Gaussian process. Such processes arise in the economic literature on learning and information acquisition—see, e.g., Bardhi, (2024), Davies, (2024), Ilut and Valchev, (2025), or Laajaj and Macours, (2025). See also Bishop, (2006, Section 6.4) or Rasmussen and Williams, (2006) for more information about Gaussian processes and their applications. Moreover, suppose the agent draws an input $x\in\mathcal{X}$ uniformly at random, and predicts the “output” $y\in\mathbb{R}$ with conditional distribution
@@
A1.2 Approximating models
Suppose the agent knows about a collection $\psi_{1},\ldots,\psi_{J}\in\mathcal{F}$ of “features” that (partially) mediate the relationship between inputs and outputs. [^44] This $J$ is the same as the depth parameter defined in Section 6. These features are linearly independent (but not necessary orthonormal) elements of the function space $\mathcal{F}$. They map inputs to known, measurable quantities. For example, if the inputs are fertilizers, then the features could map fertilizers to quantities of different nutrients.
The agent uses $\psi_{1},\ldots,\psi_{J}$ to build a “model” $m\in\mathcal{F}$ approximating the unknown function $f$. This model is a linear combination of features: there is a(n unknown) vector $\beta\equiv(\beta_{1},\ldots,\beta_{J})\in\mathbb{R}^{J}$ such that
@@
of $m(x)$ with respect to $\psi_{k}(x)$ does not depend on the input $x$. In this way, the model $m$ captures the generalizable structure of $f$ that is common to all inputs. In contrast, the model’s approximation error
@@
of $\mathcal{F}$ spanned by the features. It corresponds to a subspace
@@
of the Euclidean space $\Theta\equiv\mathbb{R}^{K}$ containing the unknown coordinate vector $\theta$. Concretely, if $\theta^{j}\equiv(\theta_{1}^{j},\ldots,\theta_{K}^{j})$ contains the (known) coordinates of the $j^{\text{th}}$ feature $\psi_{j}$ over the orthonormal basis $\mathcal{B}$, then
@@
and $1/2\leq t\leq 1/2$. Increasing $t$ from 1/2-1/2 to $1/2$ rotates $w^{(1)}$ clockwise from (1,0)(-1,0) to $(1,0)$. The value [^45] We obtain (A4) by substituting $K=2$, the prior variance matrix (8), the covariate $w^{(1)}=(\sin(\pi t),\cos(\pi t))$, and the sample size $n=1$ into (A2).
@@
If $\mathcal{S}$} is representative, then its value $\pi({\mathcal{S}})$ does not fall when $\lambda_{1},\ldots,\lambda_{K}$ undergo a mean-preserving spread.
@@
If the sample $\mathcal{S}$} is not representative, then its value can fall when the eigenvalues of $\Sigma$ undergo a MPS. This happens, for example, when ${\mathcal{S}}=\{(w^{(1)},y^{(1)})\}$ is a singleton and $w^{(1)}$ is an eigenvector of $\Sigma$ corresponding to an eigenvalue that falls under the MPS. [^46] For example, if $\theta$ has prior variance (8), then the value of ${\mathcal{S}}=\{(w^{(1)},y^{(1)})\}$ is decreasing in $\rho$ when $w^{(1)}=v_{2}$. Such a sample is “non-spanning”: the rank
@@
of $G$. The agent cannot learn about these components from $\mathcal{S}$} directly. But he can learn about them indirectly if he knows how they covary with the components that belong to $\mathrm{col}(G)$.
@@
for each $k\in\{1,\ldots,K\}$. [^47] The matrix $\operatorname{\mathbb{V}}(\theta_{\mathcal{S}})$ is invertible because it is a leading principal submatrix of an invertible matrix. Then the posterior variance matrix has trace
@@
is the prior variance of $\theta_{k}$ left unexplained $\theta_{\mathcal{S}}$. [^48] We derive (A5)–(A7) in Appendix B. The first two terms on the RHS of (A5) are “sampling errors” that depend on how much information $\mathcal{S}$} contains about the on-support components $\theta_{1},\ldots,\theta_{R}$. The third term is an “extrapolation error” that depends on how much information these components contain about the off-support components $\theta_{R+1},\ldots,\theta_{K}$. Whereas the sampling error can be reduced by collecting more (or less noisy) data on $\theta_{1},\ldots,\theta_{R}$, the extrapolation error cannot. It can only be reduced by making the on-support components of $\theta$ contain more informatio about the off-support components.
@@
This example highlights the importance of conceptual knowledge when making out-of-sample predictions. If the agent had no conceptual knowledge (and so assumed $\rho=0$), then his extrapolation error could be arbitrarily large and the value of his sample could be arbitrarily small. Having conceptual knowledge allows him to use data on the on-support components of $\theta$ to learn about the off-support components. This lowers his extrapolation error and ensures his sample has some value, even if the sampling frame is limited.
@@
    does not fall when $\lambda_{1},\ldots,\lambda_{K}$ undergo a mean-preserving spread.
@@ Appendix B Proofs
for each $k\in\{1,\ldots,K\}$ gives $\tau_{k}^{\prime}\in[\tau_{k},\tau_{k+1})$ if and only if $K/\overline{\lambda}\in[\eta_{k},\eta_{k+1})$. But $\eta_{k}$ is non-decreasing in $k$ because $\lambda_{k+1}\leq\lambda_{k}$ and therefore
@@
$\frac{\partial}{\partial\rho}\mathopen{}\mathclose{{\left[\pi^{*}\big\rvert_{\rho>\rho^{\prime}}}}\right]$ $=\frac{(K-1)\sigma^{2}}{K}\mathopen{}\mathclose{{\left(1+\mathopen{}\mathclose{{\left(\frac{1}{1+\tau\mathopen{}\mathclose{{\left(1+\rho(K-1)}}\right)}}}\right)^{2}}}\right)$
@@
    because $\lambda_{J+1}=(K-J)\lambda_{K}^{(J)}-\mathopen{}\mathclose{{\left(K-(J+1)}}\right)\lambda_{K}^{(J+1)}$. We also have
@@
    Fix $n\geq 0$. Now $\pi^{(J)}$ is non-decreasing in $J$ (by Lemma 5), so if $\pi^{(J)}\geq\pi_{0}$ then $\pi^{(J+1)}\geq\pi_{0}$. Thus
@@
$=\frac{1}{\sigma_{u}^{4}}\mathopen{}\mathclose{{\left(\Sigma^{-1}G\Sigma^{-1}+\frac{1}{\sigma_{u}^{2}}G^{2}\Sigma^{-1}+\frac{1}{\sigma_{u}^{2}}\Sigma^{-1}G^{2}+\frac{1}{\sigma_{u}^{4}}G^{3}}}\right),$
which is the sum of four matrices with strictly positive traces. Thus